Using the number 34 and 12. At some point during the lives of you and the youngest person, your age will be three times his/her age at that moment. Write an equation which models how old in years each of you will be when you are three times as old as the younger person.
I'm having a hard timing understanding this question....

Answer by rothauserc(4718)

34 + x = 3(12 +x)
34 + x = 36 +3x
2x = -2
x = -1
you will be 33 and youngest will be 11